A (field) hockey player is running north at a speed of 8 m⋅s−18 \text{ m}\cdot\text{s}^{-1}8 m⋅s−1. A ball is coming directly towards her from the north at a speed of 18 m⋅s−118\text{ m}\cdot\text{s}^{-1}18 m⋅s−1. She swings her stick forward, directly north, at a speed of 6 m⋅s−16\text{ m}\cdot\text{s}^{-1}6 m⋅s−1 measured in her own frame of reference. What is the speed of the ball with respect to the hockey stick?
1
Expert's answer
2020-07-23T08:46:10-0400

From Galileo's law of velocity addition ,Now, assume Hockey stick as point mass,thus the velocity of of the hockey stick from ground is

"v_g=8+6=14\:\text{m}\text{s}^{-1}"

Velocity of ball with respect to ground is

"v_b=-18\:\text{m}\text{s}^{-1}"

Thus,speed of the ball with respect to Hockey stick is

"v_{bs}=v_g-v_b=32\:\text{m}\text{s}^{-1}"